# PRIMT: Preference-based Reinforcement Learning with Multimodal Feedback and Trajectory Synthesis from Foundation Models

### v1

@@ 4.1 Setup
We evaluate PRIMT across a diverse set of tasks, including 2 locomotion tasks: Hopper Stand and Walker Walk from DeepMind Control (DMC) suite [44], as well as 6 articulation or rigid body manipulation tasks: Button Press, Door Open, and Sweep Into from MetaWorld suite [45], and PickSingleYCB, StackCube, and PegInsertionSide from ManiSkill suite [46]. Detailed task descriptions are provided in Appendix E. We compare PRIMT against the following baselines:
1) RL-VLM-F [14]: This baseline utilizes VLM to analyze visual renderings of trajectories to provide preference labels, representing a state-of-the-art VLM-based method; 2) RL-SaLLM-F [16]: This model employs LLM to provide preference labels based on textual descriptions of trajectories, and to modify ambiguous trajectories by generating self-improved alternatives, which are assumed to be preferred when paired with the original ones. This denotes a LLM-based method with trajectory augmentation for addressing query ambiguity; 3) PrefCLM [15]: This baseline leverages crowdsourced LLMs to provide evaluation feedback for improved feedback quality, presenting another state-of-the-art LLM-based method; 4) PrefMul: We build this baseline by directly providing the multimodal trajectory inputs used in PRIMT to a multimodal FM for evaluation, representing a naive approach to multimodal feedback; and 5) PrefGT: This baseline uses expert-designed reward functions provided by the benchmarks in a scripted teacher manner [8] to provide preference labels. This should, in theory, serve as an upper-bound oracle of PbRL performance on each task.
We also build several ablation models to assess the impact of each PRIMT component: 1) w/o.Intra: without the crowd-check mechanism and intra-modal preference fusion module; 2) w/o.Inter: without the inter-modal preference fusion module, directly selecting the modality-specific label with highest confidence as the final label; 3) w/o.ForeGen: without the foresight trajectory generation module; 4) w/o.HindAug: without the hindsight trajectory augmentation module; and 5) w/o.CauAux: without the causal auxiliary loss for counterfactuals in reward learning.
To eliminate the impact of non-model factors, we use the same trajectory inputs and CoT prompts as in PRIMT for preference label elicitation across all FM-based baselines and ablation models. The only exception is PrefCLM, which relies heavily on direct access to the environment code; for this baseline, we follow the original settings from the source paper. For all FM-based methods, we use gpt-4o as the LLM backbone. For the PbRL backbone, we use PEBBLE [7] with the uncertainty-based sampling schedule [8] for all methods, along with a consistent set of hyperparameters for the RL-based policy learning phase with SAC. This design ensures that the only difference between methods lies in the preference reward learning, allowing for a more controlled comparison.
@@ 4.2 Does PRIMT Learn Effective Rewards and Policies that Prompt Task Performance?
We first investigate whether PRIMT can learn effective reward models that lead to policies capable of solving complex tasks. Fig. 2 shows the learning curves of all methods across 8 tasks. We observe that PRIMT consistently outperforms all FM-based baselines that rely on single-modality feedback, demonstrating superior final performance and faster convergence.
However, the naive multimodal feedback method, PrefMul, performs poorly. We attribute this to the fact that directly feeding multimodal trajectory inputs to multimodal FMs without appropriate fusion can overwhelm the models, potentially even degrading performance. This highlights the need for carefully designed fusion strategies to fully leverage the complementary strengths of multimodal signals, as implemented in the hierarchical neuro-symbolic fusion strategy of PRIMT. Moreover, each component in this hierarchical design is crucial: as shown in Fig. 3, removing either intra-modal or inter-modal fusion significantly degrades task performance. This drop is especially severe when inter-modal fusion is removed, underscoring the critical role of PSL-based reasoning in PRIMT. Unlike simple rule-based fusion in w/o.Inter, PSL-based fusion more effectively captures intra-modal uncertainty, cross-modal conflicts, and the influence of trajectory context, enabling more robust preference integration from heterogeneous sources.
@@ 4.4 Does PRIMT Enhance Credit Assignment in the Reward Model?
We further investigate how the learned reward models align with the task progress at state-action level. Fig. 4 shows the normalized reward outputs from the learned reward models of PRIMT, its two ablations without full hindsight trajectory augmentation, and the baseline RL-VLM-F,
@@
along with the normalized ground-truth reward values on the same trajectories. We observe that PRIMT produces more aligned reward patterns that closely reflect task progress, while the baselines and ablations exhibit either noisy signals or high variance, indicating that their learned reward models struggle to accurately assign rewards at the state-action level, even if they capture trajectory-level preferences. This demonstrates that PRIMT, particularly its hindsight augmentation and causal auxiliary loss, enables more precise state-action-level credit assignment in reward learning.
4.5 Additional Experimental Results and Real-world Deployment
We further conducted qualitative analyses of the trajectory synthesis module and ablation studies on the influence of the FM backbone. We also demonstrated the effectiveness of PRIMT on a Kinova Jaco robot in block lifting and stacking tasks (Fig. 5). Detailed results are provided in Appendix G.
@@ 5 Conclusion and Future Work
In this work, we present PRIMT, a method that leverages foundation models for multimodal synthetic evaluation and trajectory synthesis to eliminate the need for extensive human effort and address query ambiguity and credit assignment challenges in PbRL.We demonstrated the benefits of PRIMT across a wide range of locomotion and manipulation tasks, achieving superior task performance, higher-quality synthetic feedback, and more accurately aligned reward models. While PRIMT is evaluated in single-agent robotic domains, generalizing its core components, such as multimodal evaluation and trajectory synthesis, to non-robotic and multi-agent settings represents an exciting direction for future work. Further discussion of limitations and broader impact statements are provided in Appendix H.
APPENDIX
A Limitations of Single-Modality Feedback in Robot Trajectory Evaluation .....................................................................................................................................................................................14
  • A.1 Failure Cases of VLM- and LLM-Based Preference Feedback
B Details of Keyframe Extraction .....................................................................................................................................................................................16
  • B.1 Near-Zero Velocity Detection
  • B.2 Smoothing Residual Peaks
  • B.3 Change Point Detection (CPD)
  • B.4 Combining Methods for Keyframe Extraction
C Prompts and Example Outputs .....................................................................................................................................................................................18
  • C.1 Synthetic Preference Generation
  • C.2 Foresight Trajectory Generation
  • C.3 Hindsight Trajectory Augmentation
D PSL Inference Details .....................................................................................................................................................................................20
  • D.1 Variables
  • D.2 HL-MRF Formulation
  • D.3 Łukasiewicz Relaxation
  • D.4 Template Rule Grounding
  • D.5 Example Rule Expansion
E Details on Task Environments .....................................................................................................................................................................................22
  • E.1 MetaWorld
  • E.2 ManiSkill
  • E.3 DeepMind Control Suite
F Additional Implementation Details .....................................................................................................................................................................................23
  • F.1 Baselines
  • F.2 Reward Learning
  • F.3 Policy Learning
G Additional Experimental Results .....................................................................................................................................................................................25
  • G.2 Ablation Study on FM Backbone Selection
  • G.3 Policy Visualizations of Different Methods
  • G.4 Qualitative Analysis of the Foresight Trajectory Generation Module
  • G.5 Qualitative Analysis of the Hindsight Trajectory Augmentation Module
  • G.6 Real-world Experiments
H Discussion on Limitations and Broader Impact .....................................................................................................................................................................................30
  • H.1 Assumptions and Limitations
  • H.2 Impact Statement
A.1 Failure Cases of VLM- and LLM-Based Preference Feedback
VLM Limitations and LLM Advantages
LLM Limitations and VLM Advantages
(a) VLM label accuracy generally increases as VDHigh increases.
(b) LLM label accuracy generally increases as TDHigh increases.
Appendix B Details of Keyframe Extraction
B.1 Near-Zero Velocity Detection
$\sigma=\{(s_{t},a_{t})\}_{t=1}^{T},$ (17)
$\mathbf{x}_{t}=[s_{t},a_{t}]\in\mathbb{R}^{d_{s}+d_{a}}.$ (18)
$\mathcal{K}_{\mathrm{zero}}=\{t+1\mid v_{t}<\delta_{v}\}$ (20)
B.2 Smoothing Residual Peaks
$\tilde{\mathbf{x}}_{t}=\frac{1}{2k+1}\sum_{j=-k}^{k}\mathbf{x}_{t+j}$ (22)
Frames with the top $K$ largest residual errors are selected:
B.3 Change Point Detection (CPD)
$\mathcal{K}_{\mathrm{cpd}}=\{\tau_{1},\tau_{2},\ldots,\tau_{M}\}$ (25)
B.4 Combining Methods for Keyframe Extraction
Appendix C Prompts and Example Outputs
C.1 Synthetic Preference Generation
Table 1: Task description for each task used in our experiments.
Task Name Task Description
Button Press to press a button on a surface
Door Open to open a door with a revolving joint.
Sweep Into to sweep the object into the target area
(a) Illustration of the VLM prompt.
(b) Illustration of the LLM prompt.
(a) Foresight trajectory generation prompt.
(b) Example inputs on the Metaworld environment.
C.2 Foresight Trajectory Generation
C.3 Hindsight Trajectory Augmentation
Appendix D PSL Inference Details
Variables.
HL-MRF Formulation.
Łukasiewicz Relaxation.
$A_{1}\widetilde{\land}A_{2}$ $=\max\{0,A_{1}+A_{2}-1\}$ (29)
$A_{1}\widetilde{\lor}A_{2}$ $=\min\{A_{1}+A_{2},1\}$ (30)
$\widetilde{\lnot}A_{1}$ $=1-A_{1}$ (31)
Template Rule Grounding.
Example Rule Expansion.
Consider the rule:
Its logical equivalent is:
and the resulting hinge-loss potential is:
Appendix E Details on Task Environments
E.1 MetaWorld
Button Press
(a) Button Press Task
(b) Door Open Task
(c) Sweep Into Task
Figure 13: Visualizations of MetaWorld Tasks
Door Open
Sweep Into
E.2 ManiSkill
PickSingleYCB
(a) PickSingleYCB Task
(b) StackCube Task
(c) PegInsertionSide Task
Figure 14: Visualizations of ManiSkill Tasks
StackCube
PegInsertionSide
E.3 DeepMind Control Suite
Hopper Stand
(a) Hopper Stand Task
(b) Walker Walk Task
Figure 15: Visualizations of DeepMind Control Suite Tasks
Walker Walk
Appendix F Additional Implementation Details
F.1 Baselines
We implemented the baselines RL-VLM-F [14], RL-SaLLM-F [16], and PrefCLM [15] using the source code released by the authors. To eliminate non-model differences, we replaced the original prompts in RL-VLM-F and RL-SaLLM-F with the same three-step chain-of-thought prompt used in PRIMT. This three-step prompt can be considered an enhanced version of the two-step reasoning used in their original papers, with an additional self-verification step. The trajectory inputs are also the same as in PRIMT: $kvis(\sigma)$ for RL-VLM-F and $text(\sigma)$ for RL-SaLLM-F.
The PrefGT baseline follows the scripted teacher approach introduced in [8]. It serves as an upper-bound oracle that has complete access to the benchmark’s ground-truth reward function. Although such access is infeasible in real-world robotic systems, it provides a useful reference to assess the best possible performance that any preference-based learning method could achieve.
$R(\sigma)=\sum_{t=1}^{T}r(s_{t},a_{t})$
F.2 Reward Learning
We adopt PEBBLE [7] as the PbRL backbone for all methods. PEBBLE first performs unsupervised pre-training to maximize state entropy and initialize the policy, followed by off-policy reinforcement learning using Soft Actor-Critic (SAC) for policy optimization. During training, all reward values in the replay buffer are relabeled whenever a new reward model is learned.
For the reward model, we implement the image-based reward network as described in their original paper for RL-VLM-F. For all other baselines using the standard PEBBLE [7] reward model, we adopt a 3-layer ensemble architecture following their design.
For selecting informative queries, we adopt an uncertainty-based query selection strategy following [8]. Specifically, we measure the uncertainty of preference predictions either by computing the variance across an ensemble of preference predictors. We then select the top-$N_{\text{query}}$ trajectory segment pairs with the highest uncertainty as query candidates to elicit preference labels.
Table 2: Hyperparameters used for reward learning.
Hyperparameter Value
Trajectory segment length 100
Feedback frequency 5000
Maximum feedback samples 20000
Number of foresight trajectories 200
Max counterfactuals per trajectory 5
F.3 Policy Learning
For all methods evaluated in this work, we follow PEBBLE [7] and adopt Soft Actor-Critic (SAC) as the off-policy reinforcement learning algorithm. To ensure fair comparison, all methods share the same actor-critic architecture and hyperparameter settings during policy learning.
Table 3: Hyperparameters used for SAC.
Hyperparameter Value Hyperparameter Value
Initial temperature 0.1 Batch Size 1024
Learning rate 0.0003 Optimizer Adam
Critic target update freq 2 Critic EMA $\tau$ 0.005
$(\beta_{1},\beta_{2})$ (0.9, 0.999) Discount $\gamma$ 0.99
Hidden units per each layer 1024
Appendix G Additional Experimental Results
  • Policy visualizations comparing PRIMT and baseline behaviors.
  • Real-world deployment on a Kinova Jaco robot.
G.2 Ablation Study on FM Backbone Selection
Figure 18: Ablation study on FM backbone selection.
G.5 Policy Visualizations of Different Methods
G.6 Real-world Experiments
We further evaluate our method in the real world on two tasks from the Robosuite benchmark [64]: Block Lifting and Block Stacking. In the Block Lifting task, the goal is to control a robotic arm to lift a cube to a certain height. In the Block Stacking task, the objective is to place one cube on top of another. We adopt PrefCLM [15] as a real-world baseline.
Policies are first trained in simulation using Robosuite [64], and then deployed on a physical Kinova Jaco2 robotic arm. Performance results are shown in Figure 23. PRIMT demonstrates superior performance on both tasks, highlighting its effectiveness in the real world. More videos of real-world experiments are on our project website.
Appendix H Discussion on Limitations and Broader Impact
H.1 Assumptions and Limitations
In this work, we make the following assumptions, each grounded in prior work and empirical observations: 1) We assume that foundation models (FMs), including LLMs and VLMs, are pretrained on large-scale and diverse corpora, enabling generalization across a wide range of environments and tasks; 2) We assume that VLMs are capable of analyzing sequences of visual inputs to extract spatial and semantic cues, while LLMs can interpret temporal structures, reason over structured inputs, and generate coherent task-related feedback; 3) We assume that LLMs, when guided by structured prompts and domain-aligned vocabulary, are capable of generating or editing robot trajectories in a coherent and task-relevant manner. This assumption builds upon recent work demonstrating the ability of LLMs to plan actions [41], generate code-based controllers [38], and perform trajectory synthesis via natural language [36]; and 4) We assume that task progress can be inferred at least partially through either visual sequences (e.g., keyframe differences, spatial transitions) or textual descriptions (e.g., arrays of actions or states).
Our work is fundamentally built upon FMs, and therefore, like other FM-based methods, it inherits certain dependencies on the capabilities and limitations of these models. While this reliance is inherent to the design, we have taken steps to mitigate it: for example, leveraging multimodal feedback to enhance the quality of preference signals and adopting a structured code-generation paradigm to improve the quality of generated trajectories. As shown in our ablation study on LLM backbone selection, PRIMT maintains reasonable performance even when using a weaker model (gpt-4o-mini), indicating robustness to FM capacity. Another limitation lies in the cost associated with FM usage. Specifically, our multimodal feedback generation, fusion module, and bidirectional trajectory synthesis involve multiple FM calls. That said, the trend toward cheaper APIs and more accessible open-source FMs partially offsets this concern, and future improvements in local FM efficiency may further alleviate the issue. Lastly, although we evaluate PRIMT across diverse tasks, it is still limited to robotic manipulation and locomotion domains. Generalizing to other domains (e.g., multi-agent planning, open-world exploration) may require adapting the feedback prompt structures, reward modeling objectives, or modality alignment strategies.
H.2 Impact Statement
This work explores the integration of foundation models into preference-based reinforcement learning (PbRL), aiming to improve learning efficiency and robustness through multimodal feedback and trajectory synthesis. By leveraging VLMs and LLMs, our approach reduces the need for extensive human supervision, potentially broadening access to PbRL techniques in domains with limited annotation resources.

### v2

@@ 4.1 Setup
We evaluate PRIMT across a diverse set of tasks, including 2 locomotion tasks: Hopper Stand and Walker Walk from DeepMind Control (DMC) suite [tassa2018deepmind], as well as 6 articulation or rigid body manipulation tasks: Button Press, Door Open, and Sweep Into from MetaWorld suite [yu2020meta], and PickSingleYCB, StackCube, and PegInsertionSide from ManiSkill suite [gu2023maniskill2]. Detailed task descriptions are provided in Appendix E. We compare PRIMT against the following baselines:
  • RL-VLM-F [wang2024rl]: This baseline utilizes VLM to analyze visual renderings of trajectories to provide preference labels, representing a state-of-the-art VLM-based method.
  • RL-SaLLM-F [tu2024online]: This model employs LLM to provide preference labels based on textual descriptions of trajectories, and to modify ambiguous trajectories by generating self-improved alternatives, which are assumed to be preferred when paired with the original ones. This denotes a LLM-based method with trajectory augmentation for addressing query ambiguity.
  • PrefCLM [wang2025prefclm]: This baseline leverages crowdsourced LLMs to provide evaluation feedback for improved feedback quality, presenting another state-of-the-art LLM-based method.
  • PrefMul: We build this baseline by directly providing the multimodal trajectory inputs used in PRIMT to a multimodal FM for evaluation, representing a naive approach to multimodal feedback.
  • PrefGT: This baseline uses expert-designed reward functions provided by the benchmarks in a scripted teacher manner [lee2021b] to provide preference labels. This should, in theory, serve as an upper-bound oracle of PbRL performance on each task.
We also build several ablation models to assess the impact of each PRIMT component:
  • w/o.Intra: without the crowd-check mechanism and intra-modal preference fusion module.
  • w/o.Inter: without the inter-modal preference fusion module, directly selecting the modality-specific label with highest confidence as the final label.
  • w/o.ForeGen: without the foresight trajectory generation module.
  • w/o.HindAug: without the hindsight trajectory augmentation module.
  • w/o.CauAux: without the causal auxiliary loss for counterfactuals in reward learning.
To eliminate the impact of non-model factors, we use the same trajectory inputs and CoT prompts as in PRIMT for preference label elicitation across all FM-based baselines and ablation models. The only exception is PrefCLM, which relies heavily on direct access to the environment code; for this baseline, we follow the original settings from the source paper. For all FM-based methods, we use gpt-4o as the LLM backbone. For the PbRL backbone, we use PEBBLE [lee2021pebble] with the uncertainty-based sampling schedule [lee2021b] for all methods, along with a consistent set of hyperparameters for the RL-based policy learning phase with SAC. This design ensures that the only difference between methods lies in the preference reward learning, allowing for a more controlled comparison.
@@ 4.2 Does PRIMT Learn Effective Rewards and Policies that Prompt Task Performance?
We first investigate whether PRIMT can learn effective reward models that lead to policies capable of solving complex tasks. Fig. 2 shows the learning curves of all methods across 8 tasks. We observe that PRIMT consistently outperforms all FM-based baselines that rely on single-modality feedback, demonstrating superior final performance and faster convergence. However, the naive multimodal feedback method, PrefMul, performs poorly. We attribute this to the fact that directly feeding multimodal trajectory inputs to multimodal FMs without appropriate fusion can overwhelm the models, potentially even degrading performance. This highlights the need for carefully designed fusion strategies to fully leverage the complementary strengths of multimodal signals, as implemented in the hierarchical neuro-symbolic fusion strategy of PRIMT. Moreover, each component in this hierarchical design is crucial: as shown in Fig. 3, removing either intra-modal or inter-modal fusion significantly degrades task performance. This drop is especially severe when inter-modal fusion is removed, underscoring the critical role of PSL-based reasoning in PRIMT. Unlike simple rule-based fusion in w/o.Inter, PSL-based fusion more effectively captures intra-modal uncertainty, cross-modal conflicts, and the influence of trajectory context, enabling more robust preference integration from heterogeneous sources.
@@ 4.4 Does PRIMT Enhance Credit Assignment in the Reward Model?
We further investigate how the learned reward models align with the task progress at state-action level. Fig. 4 shows the normalized reward outputs from the learned reward models of PRIMT, its two ablations without full hindsight trajectory augmentation, and the baseline RL-VLM-F, along with the normalized ground-truth reward values on the same trajectories. We observe that PRIMT produces more aligned reward patterns that closely reflect task progress, while the baselines and ablations exhibit either noisy signals or high variance, indicating that their learned reward models struggle to accurately assign rewards at the state-action level, even if they capture trajectory-level preferences. This demonstrates that PRIMT, particularly its hindsight augmentation and causal auxiliary loss, enables more precise state-action-level credit assignment in reward learning. Extra quantitative results using the $R^{2}$ coefficient are provided in Appendix G.8.
4.5 Additional Experimental Results
We conducted additional analyses to provide further validation insights of the proposed framework. These include: qualitative evaluations of the trajectory synthesis module (Appendices G.3 and G.4), visualization of policy outcomes across different methods (Appendix G.5), comparison with the dense-reward RL baseline (Appendix G.7), and preliminary experiments on a more complex dual-arm manipulation task (Appendix G.6).
We also performed ablation studies on the influence of the foundation model backbone (Appendix G.2). The results show good potential for more accessible deployment with smaller or cheaper foundation models: with GPT-4o-mini, we achieved a 94% reduction in cost with a 16–26% drop in performance, leading to a 13× improvement in cost–performance efficiency.
4.6 Real-world Deployment
We further demonstrated the effectiveness of PRIMT on a Kinova Jaco robot in block lifting and stacking tasks (Fig. 5). Detailed experimental settings and results are provided in Appendix G.9.
@@ 5 Conclusion and Future Work
In this work, we presented PRIMT, a method that leverages foundation models for multimodal synthetic evaluation and trajectory synthesis to reduce human effort and address query ambiguity and credit-assignment challenges in preference-based reinforcement learning (PbRL). We demonstrated the advantages of PRIMT across a wide range of locomotion and manipulation tasks, achieving superior task performance, higher-quality synthetic feedback, and more accurately aligned reward models.
A limitation of this study involves the increased cost associated with foundation-model usage due to the multimodal components of the framework. To provide a clearer understanding of the resource requirements, we conducted a detailed analysis of computational usage and the corresponding cost–performance trade-offs, presented in Appendix H.1. We observe that PRIMT strikes a good balance between performance and cost-effectiveness, providing a practical path toward scalable preference learning.
Looking forward, while PRIMT was evaluated in single-agent robotic domains, extending its key components, such as multimodal evaluation and trajectory synthesis, to non-robotic and multi-agent settings represents an exciting direction for future work. Further discussion of assumptions, limitations, and broader impacts is provided in Appendix H.
Acknowledgments
The authors would like to thank Jinjin Cai from Purdue University for insightful comments on the figures and visualizations. This work was partially supported by the Ministry of Science and ICT (MSIT), Korea, under the Global Research Support Program in the Digital Field (RS-2024-00425354), supervised by the Institute for Information & Communications Technology Planning & Evaluation (IITP). The support applies to authors Hong and Min.